OpenStudy (anonymous):

which expression defines the given series for seven terms? -3+(-1)+1+...

OpenStudy (anonymous):

It seems like the series is: -3 + 2n. So, 7th term = -3 + 2*7 = -3 + 14 = 11

OpenStudy (anonymous):

I am assuming first term is 0th term (n=0).